OpenStudy (anonymous):

Evaluate the integral by making the given substitution. the integral of cos^6 θ sin θ dθ, u = cos θ

zepdrix (zepdrix):

So with a substitution, we want to replace ALL of the Theta's and D(Theta) with U's and du. Let \[\large u=\cos \theta\]Taking the derivative gives us,\[\large du=-\sin \theta d \theta\] But our equation includes a sin theta, without the negative, so let's move the negative to the other side of that equation. \[\large -du= \sin \theta d \theta\]

zepdrix (zepdrix):

So now we can replace the cos theta with something involving u, and we can replace the sin theta d(theta) with something involving du, and it should simplify our integral.

zepdrix (zepdrix):

\[\large \int\limits\limits (\cos \theta)^6 (\sin \theta d \theta) \rightarrow \quad \int\limits\limits (u)^6(-du)\]

zepdrix (zepdrix):

Understand what we did there? :D Substitutions can be a little tricky to get used to. It's not as easy as differentiation was ^^

OpenStudy (anonymous):

ohh okay i see what i did wrong now. thank you!